# Towards Calibrated Robust Fine-Tuning of Vision-Language Models

### v5

@@ 4.1 Robust fine-tuning with constrained multimodal contrastive learning
In order to adapt a pre-trained VLM on image classification tasks, the cross-entropy loss is the most common choice as an objective function. However, there are emerging shreds of evidence supporting the use of contrastive loss (CL) for robust adaptation [61, 43, 13], especially when the model is pre-trained via CL. Witnessing its empirical success on OOD generalization [13], we leverage a CL-based learning strategy for VLM fine-tuning. CLIP consists of image encoder $f_{\theta_{v}}(\cdot)=f_{\hat{\theta}_{v}}(\cdot)W_{v}^{T}$ and text encoder $g_{\theta_{l}}(\cdot)=g_{\hat{\theta}_{l}}(\cdot)W_{l}^{T}$, where the encoders are composed with backbone models $(f_{\hat{\theta}_{v}}(\cdot),g_{\hat{\theta}_{l}}(\cdot))$ and projection layers $(W_{v}^{T},W_{l}^{T})$. The encoders produce $L_{2}$-normalized representations to compute the similarity between image and text inputs. Given $N$ pairs of (image, text) $\{(I_{i},T_{i})\}_{i=1}^{N}$, a common form of multimodal contrastive loss (MCL) can be written as
@@
where $\theta=\{\theta_{v},\theta_{l}\}$ are the parameters of image and text encoders and $R(\theta_{v},\theta_{l})$ reflects general regularization strategies in CL [6, 17]. We update both image and text encoders during fine-tuning as done in the pre-train phase and use OpenAI templates [40] to create (image, text) pairs from a downstream classification dataset that consists of (image, class) pairs.
@@
$\mathcal{L}_{\text{MCL-con}}(\theta):=\mathcal{L}_{\text{MCL}}(\theta)+\lambda_{\text{OC}}\mathcal{L}_{\text{OC}}(W_{v}),\quad\mathcal{L}_{\text{OC}}(W_{v})=||W_{v}W_{v}^{T}-\mathbf{I}||_{F}^{2},$ (4)
where $\mathbf{I}$ is an identity matrix that has the same shape with $W_{v}$ and $\lambda_{\text{OC}}$ is a strength of the orthogonality constraint [^1] While it is also possible to inject constraint on the text projection matrix $W_{l}$, we only do it for $W_{v}$ because our concern is about the singular value of image representations for downstream tasks. We observed degradations of accuracy (ID:$-0.01$, OOD:$-0.15$) and ECE (ID:$-0.002$, OOD:$-0.02$) by adding the constraint on $W_{v}$.. While recklessly increasing the singular values might hinder ID adaptation, our orthogonal constraint mitigates the degradation of performance by pursuing not only the smallest singular values to be large but also the largest singular values to be small which is important for generalization on ID data [60]. Interestingly, this contrastive loss with regularization terms can be viewed as a singular value decomposition (SVD) on the cross-covariance matrix of image-text representations with an orthogonality constraint.
@@
$\mathcal{\operatorname*{arg\,min}}_{W_{v},W_{l}}\mathcal{L}_{\text{MCL-con}}(W)$ $:={1\over N}\sum_{i=1}^{N}-\log\frac{\exp(W_{v}\hat{I}_{i}\cdot W_{l}\hat{T}_{i})}{\sum_{j=1}^{N}\exp(W_{v}\hat{I}_{i}\cdot W_{l}\hat{T}_{j})}$
$+{1\over N}\sum_{i=1}^{N}-\log\frac{\exp(W_{v}\hat{I}_{i}\cdot W_{l}\hat{T}_{i})}{\sum_{j=1}^{N}\exp({W_{v}\hat{I}_{j}\cdot W_{l}\hat{T}_{i})}}+R(W_{v},W_{l})+\lambda_{\text{OC}}||W_{v}W_{v}^{T}-\mathbf{I}||_{F}^{2}$
$\approx\quad\operatorname*{arg\,max}_{W_{v},W_{l}}\text{SVD}(S(\beta))$ $:=\;\mathrm{tr}(W_{v}S(\beta){W_{l}}^{T})-(\rho/2)||W_{v}^{T}W_{l}||_{F}^{2}\quad\text{subject to}\;||W_{v}W_{v}^{T}-\mathbf{I}||_{F}^{2}=0,$
where we adopt $R(W_{v},W_{l})=(\rho/2)||W_{v}^{T}W_{l}||_{F}^{2}$ for $\rho>0$ as a regularization term to promote the encoders to capture diverse features as in Ji et al. [21]. Here, we assume that the input of $\mathcal{L}_{\text{MCL-con}}$ is the penultimate representation of VLM’s encoders, i.e., $(\hat{I}_{i},\hat{T}_{i})=(f_{\hat{\theta}_{v}}(I_{i}),g_{\hat{\theta_{l}}}(T_{i}))$, and $W=\{W_{v},W_{l}\}$ is a set of projection matrices (i.e., last layers of each encoder). This equivalence between $\mathcal{L}_{\text{MCL}}$ and SVD allows us to understand the working mechanism of our proposed objective. That is, minimizing $\mathcal{L}_{\text{MCL-con}}$ can be interpreted as finding a good approximation over the cross-modal covariance matrix by seeking the direction of maximal variation among image-text representations, while enforcing a large smallest singular value of the covariance matrix of image representation via orthogonality constraint on $W_{v}$. In §5, we validate that $\mathcal{L}_{\text{MCL-con}}$ significantly increases the smallest singular values and results in better OOD generalization and calibration on downstream tasks.
@@ 4.2 Calibration during robust fine-tuning
In the previous section, we devise a new multimodal contrastive loss that promotes large $\sigma_{min}(\tilde{\Sigma}_{\mathcal{D}_{\text{ID}}})$. We now address the next component standing for ID calibration, which is another crucial component according to our theoretical analysis. While there are numerous approaches to enhance calibration during neural network training [32, 48, 45, 1], we notice the promising results of knowledge distillation-based calibration approaches [55, 62]. These approaches encourage the model to learn from input-dependent smoothed labels that effectively mitigate the overconfidence issue, which is commonly associated with miscalibration. We employ a self-distillation (SD) method for ID calibration in that distilling supervision would help avoid overconfidence.
@@
Specifically, we first initialize both teacher and student networks with a pre-trained CLIP model (including both image and text encoders); update the student model using gradient descent for every iteration while slowly updating the teacher model that has $\psi=\{\psi_{v},\psi_{l}\}$ as parameters using EMA with the momentum of $\alpha$ at every $t>1$ iteration, i.e., $\psi\leftarrow\alpha\psi+(1-\alpha)\theta$. Rather than hosting another VLM or pre-trained CLIP as a teacher model, we adopt a self-evolving EMA network as a teacher, in accordance with the notable observation of the impressive robustness of weight-space ensembling among identical structure networks [53, 41]. With the EMA teacher $\psi$ and the learning student $\theta$, we construct a self-distillation loss term for $N$ data pairs as:
$\mathcal{L}_{\text{SD}}(\theta):=\sum_{i=1}^{N}[KL(\tilde{q}^{I}_{i}||q^{I}_{i})+KL(\tilde{q}^{T}_{i}||q^{T}_{i})],$ (5)
@@
where $KL$ denotes Kullback–Leibler divergence, $q^{I}_{i}=\texttt{softmax}(\{f_{\theta_{v}}(I_{i})\cdot g_{\theta_{l}}(T_{j})\}_{j=1}^{N})$ and $q^{T}_{i}=\texttt{softmax}(\{f_{\theta_{v}}(I_{j})\cdot g_{\theta_{l}}(T_{i})\}_{j=1}^{N})$ are student outputs, and $\tilde{q}^{I}_{j}$ and $\tilde{q}^{T}_{j}$ are teacher outputs which are similarly defined by replacing the student parameter $\theta$ with teacher $\psi$. Presumably, label smoothing (LS) [47] achieves effects similar to what we intended, but we argue that LS would have overconfidence issues, as demonstrated in the Appendix §B.
@@
We complete the learning objective as a summation of $\mathcal{L}_{\text{MCL-con}}$ and $\mathcal{L}_{\text{SD}}$ with a coefficient $\lambda_{\text{SD}}$, i.e., $\mathcal{L}=\mathcal{L}_{\text{MCL}}+\lambda_{\text{OC}}\mathcal{L}_{\text{OC}}+\lambda_{\text{SD}}\mathcal{L}_{\text{SD}}$. The novel combination of these two components contributes to ensuring a larger smallest singular value of image representation and ID calibration simultaneously, which induces calibrated robust fine-tuning (CaRot) on diverse distribution shifts.
@@ 5 Experiments
In § 5.1, we first show empirical evidence of the error bounds that we derived in §3. We then provide experimental setup (§5.2) and discuss the empirical analyses on benchmark datasets (§5.3).
@@ 5.1 Numerical analysis on error bounds
Our theoretical analysis in §3 revealed the capability of managing OOD classification and calibration errors using the same quantities over ID domain at the same time. Before conducting real-world evaluations, we verify our theoretical analysis with our toy experiments that simulate distribution shifts. To be specific, we generate a binary classification dataset with 1000-dimensional Gaussian features where the mean of features are partly shifted across different test environments (ID, OOD). We train a three-layer network with regularization terms: Born-Again-Network (BAN)-style self-distillation [10] for calibration ($\mathcal{L}_{\text{SD}}$) and the orthogonal constraint (shown in eq.(4)) for the singular value ($\mathcal{L}_{\text{OC}}$). Detailed descriptions of the experimental setup are provided in Appendix §A.1.
Figure 3: Analysis of error bounds on synthetic data. Plots on the left side show RHS (x-axis) and LHS (y-axis; MSE for eq.(2) and ECE for eq.(1)) of the inequalities in §3. We denote MSE for the mean squared error, $\mathcal{L}_{\text{OC}}$ for the singular value regularization, and $\mathcal{L}_{\text{SD}}$ for the calibration regularization.
@@
Figure 3 visualizes the results of Pearson correlation analysis between the average of $1/\sigma_{min}(\tilde{\Sigma}_{\mathcal{D}_{\text{ID}}})$ and ECE from ID samples and OOD MSE/ECE over 111 trained models. Here, we observe strong correlations between the average of $1/\sigma_{min}(\tilde{\Sigma}_{\mathcal{D}_{\text{ID}}})$ and ID ECE (x-axis), and OOD classification and calibration errors (y-axis). Additional results on the best models per each regularization term are showcased on the Table 5.1, which also indicates that reducing the upper bound results in better OOD generalization and calibration. These analyses demonstrate our upper bound eq.(3.1) empirically holds.
@@ 5.2 Evaluation on distribution shift benchmarks
Benchmark datasets. We consider IN-V2 [42], IN-R [18], IN-A [19], IN-S [52], and ObjectNet [2]. We refer to these five datasets as OOD throughout the following sections unless it is specified as a different dataset.
Baseline methods. We benchmark CaRot alongside zero-shot inference (ZS) and fine-tuning methods: standard fine-tuning (FT), LP-FT [25], FLYP [13], and Lipsum-FT [36]. Refer §A for further details of each method. In Appendix Table A, B, and C, we compare results when applying post-hoc techniques for robust fine-tuning (weight ensembling; WiSE-FT [53]) and post-hoc calibration (temperature scaling; TS [14]) methods.
@@
Evaluation results. Table 2 and 3 highlight our argument that CaRot significantly enhances both generalization and calibration on OOD data. Under distribution shifts from IN to -V2, -R, -A, -S, and ObjectNet, CaRot favorably compares with the existing best fine-tuning methods by margin of 1.51 and 0.0123 for OOD top-1 accuracy and ECE, respectively, averaged over five shifted datasets. The reliability diagrams are provided in Appendix Figure B. We report the results with different VLM backbones, RN50 and ViT-L/14, in Table 7 (See Table G and H for details). CaRot consistently outperforms the baseline methods for these backbones, too. Furthermore, in Table 6, we provide additional comparisons with CAR-FT  [28], Model Stock [20] and ARF [16]. [^3] We report the averaged accuracy over four shifted datasets in this Table 6. Results for methods marked with an asterisk (*) are referenced from the respective original papers. The ECE results are not included due to the missing evaluations from the original papers.
5.3 Empirical studies
Ablation study. In Table 4, we provide ablation study results to show the impacts of each component of CaRot. In line with our hypothesis, the results confirm that all three components boost OOD accuracy and calibration performance. The comparison of adopting and not adopting $\mathcal{L}_{\text{MCL}}$ (we followed the naive fine-tuning approach for the latter) ascertains that employing contrastive loss as a fine-tuning objective is superior to cross-entropy loss for ID/OOD accuracy, consistent with the previous observations [13], and even extends to improvements in calibration as well. The ablations of the orthogonality constraint and adopting self-distillation validate our rationale behind adding the terms to our learning objective, where we expect them to lower the upper bound of OOD classification and calibration errors. Together, constraining the singular values of image representation on MCL and distilling EMA teacher’s predictions show the best results which aligned with results from the demonstration of error bounds in Fig 3. We further analyze the impact of each component by varying the strength coefficients in Table 5. We observe that the increased intensity of constraint improves ID and OOD performance to some degree, but there is a slight decline when the intensity is too strong. Meanwhile, the strength of self-distillation positively correlated with OOD accuracy and ECE, but there is a negative correlation between ID accuracy and ECE, which reflects the inevitable trade-off between ID adaptation and OOD generalization (Table F and E in Appendix provide further details).

### v7

@@ 4.1 Robust fine-tuning with constrained multimodal contrastive learning
To adapt a pre-trained VLM on image classification tasks, the cross-entropy loss is the most common choice as an objective function. However, there are emerging shreds of evidence supporting the use of contrastive loss (CL) for robust adaptation [70, 50, 17], especially when the model is pre-trained via CL. Witnessing its empirical success on OOD generalization [17], we leverage a CL-based learning strategy for VLM fine-tuning. CLIP consists of an image encoder $f_{\theta_{v}}(\cdot)=f_{\hat{\theta}_{v}}(\cdot)W_{v}$ and a text encoder $g_{\theta_{l}}(\cdot)=g_{\hat{\theta}_{l}}(\cdot)W_{l}$, where encoders are composed with backbone models $(f_{\hat{\theta}_{v}}(\cdot),g_{\hat{\theta}_{l}}(\cdot))$ and projection matrices $(W_{v}\in\mathbb{R}^{d_{v}\times r},W_{l}\in\mathbb{R}^{d_{l}\times r})$. The encoders produce $L_{2}$-normalized representations to compute the similarity between image and text inputs. Given $N$ pairs of (image, text) $\{(I_{i},T_{i})\}_{i=1}^{N}$, a common form of multimodal contrastive loss (MCL) can be written as
@@
where $\theta=\{\theta_{v},\theta_{l}\}$ are the parameters of image and text encoders and $R(\theta_{v},\theta_{l})$ reflects a general regularization strategy in CL [7, 21]. We update both image and text encoders during fine-tuning as done in the pre-train phase and use OpenAI templates [47] to create (image, text) pairs from a downstream classification dataset that consists of (image, class) pairs.
Figure 2: Overview of CaRot. We fine-tune a VLM using a multimodal contrastive loss with an orthogonality constraint on visual projection layer (eq.(4)) and self-distillation $\mathcal{L}_{\text{SD}}$ (eq.(5)) that takes predictions of EMA teacher $\psi$ as soft target labels to train the student model $\theta$. The darker and the lighter elements denote values closer to 1 and 0, respectively. Both teacher and student models share identical VLM architecture consisting of image $f_{\theta_{v}}:=[f_{\hat{\theta}_{v}};W_{v}]$ and text $g_{\theta_{l}}:=[g_{\hat{\theta}_{l}};W_{l}]$ encoders, where $W$ is the last projection layer. Given (image, text) pair data, the model outputs the pair-wise similarity score for in-batch image-text representations.
@@
$\mathcal{L}_{\text{MCL-con}}(\theta):=\mathcal{L}_{\text{MCL}}(\theta)+\lambda_{\text{OC}}\mathcal{L}_{\text{OC}}(W_{v}),\quad\mathcal{L}_{\text{OC}}(W_{v})=||W_{v}^{T}W_{v}-\mathbf{I}||_{F}^{2},$ (4)
where $\mathbf{I}$ is an identity matrix that has the same shape with $W_{v}^{T}W_{v}$ and $\lambda_{\text{OC}}$ is a strength of the orthogonality constraint [^1] While it is also possible to inject constraint on the text projection matrix $W_{l}$, we only do it for $W_{v}$ because our concern is about the singular value of image representations for downstream tasks. We observed degradations of accuracy (ID:$-0.01$, OOD:$-0.15$) and ECE (ID:$-0.002$, OOD:$-0.02$) by adding the constraint on $W_{l}$.. While recklessly increasing the singular values might hinder ID adaptation, our orthogonal constraint mitigates the degradation of performance by pursuing not only the smallest singular values to be large but also the largest singular values to be small which is important for generalization on ID data [69]. Interestingly, this contrastive loss with regularization terms can be viewed as a constrained singular value decomposition (SVD) with a cross-covariance matrix of image-text representations where the orthogonality constraint is applied.
@@
$\operatorname*{arg\,min}_{W_{v},W_{l}}\mathcal{L}_{\text{MCL-con}}(W)$ $:={1\over 2N}\sum_{i=1}^{N}-\log\frac{\exp(W_{v}\hat{I}_{i}\cdot W_{l}\hat{T}_{i})}{\sum_{j=1}^{N}\exp(W_{v}\hat{I}_{i}\cdot W_{l}\hat{T}_{j})}$
$+{1\over 2N}\sum_{i=1}^{N}-\log\frac{\exp(W_{v}\hat{I}_{i}\cdot W_{l}\hat{T}_{i})}{\sum_{j=1}^{N}\exp({W_{v}\hat{I}_{j}\cdot W_{l}\hat{T}_{i})}}+R(W_{v},W_{l})+\lambda_{\text{OC}}||W_{v}^{T}W_{v}-\mathbf{I}||_{F}^{2}$
$\approx\quad\operatorname*{arg\,max}_{W_{v},W_{l}}\text{SVD}(S(\beta))$ $:=\;\mathrm{tr}(W_{v}^{T}S(\beta){W_{l}})-(\rho/2)||W_{v}W_{l}^{T}||_{F}^{2}\quad\text{subject to}\;||W_{v}^{T}W_{v}-\mathbf{I}||_{F}^{2}=0,$
where we adopt $R(W_{v},W_{l})=(\rho/2)||W_{v}W_{l}^{T}||_{F}^{2}$ for $\rho>0$ as a regularization term to promote the encoders to capture diverse features as in Ji et al. [27]. Here, we assume that the input of $\mathcal{L}_{\text{MCL-con}}$ is the penultimate representation of VLM’s encoders, i.e., $(\hat{I}_{i},\hat{T}_{i})=(f_{\hat{\theta}_{v}}(I_{i}),g_{\hat{\theta_{l}}}(T_{i}))$, and $W=\{W_{v},W_{l}\}$ is a set of projection matrices (i.e., last layers of each encoder). This connection between $\mathcal{L}_{\text{MCL}}$ and SVD allows us to understand the working mechanism of our proposed objective. That is, minimizing $\mathcal{L}_{\text{MCL-con}}$ can be interpreted as finding a good rank-$r$ (dimensionality of the image-text projection space) approximation of the cross-modal covariance matrix by seeking the direction of large co-variation among image-text representations, while the solution space is constrained by enforcing an orthogonality condition on the collection of vision-side singular vectors $W_{v}$ to achieve the larger effective rank of both the projection matrix $W_{v}$ and the image representation matrix (See Appendix §B for further explanation on the effective rank and the smallest singular value). In §5, we validate that $\mathcal{L}_{\text{MCL-con}}$ significantly increases the smallest singular values and results in better OOD generalization and calibration on downstream tasks.
@@ 4.2 Calibration during robust fine-tuning
In the previous section, we devise a new multimodal contrastive loss that promotes large $\sigma_{min}(\tilde{\Sigma}_{\mathcal{D}_{\text{ID}}})$. We now address the next component standing for ID calibration, which is another crucial component according to our theoretical analysis. While there are numerous approaches to enhance calibration during neural network training [38, 56, 52, 1], we notice the promising results of knowledge distillation-based calibration approaches [64, 71]. These approaches encourage the model to learn from input-dependent smoothed labels that effectively mitigate the overconfidence issue, which is commonly associated with miscalibration. Therefore, we employ a self-distillation (SD) method for ID calibration in that distilling the similarity score map would help avoid overconfidence.
@@
Specifically, we first initialize both teacher and student networks with a pre-trained CLIP model (including both image and text encoders); update the student model using gradient descent for every iteration while slowly updating the teacher model that has $\psi=\{\psi_{v},\psi_{l}\}$ as parameters using EMA with the momentum of $\alpha$ at every $t>1$ iteration, i.e., $\psi\leftarrow\alpha\psi+(1-\alpha)\theta$. Rather than hosting another VLM or fixed pre-trained CLIP as a teacher model, we adopt a self-evolving EMA network as a teacher, observing its successful usage on the weight-space ensemble between homogeneous models [61, 48], robust self-supervised learning methods [2, 45], as well as regularization [71]. With the EMA teacher $\{f_{\psi_{v}}(\cdot),g_{\psi_{l}}(\cdot)\}$ and the learning student $\{f_{\theta_{v}}(\cdot),g_{\theta_{l}}(\cdot)\}$, we construct a self-distillation loss term for $N$ data pairs as:
$\mathcal{L}_{\text{SD}}(\theta):=\frac{1}{N}\sum_{i=1}^{N}[KL(\tilde{q}^{I}_{i}||q^{I}_{i})+KL(\tilde{q}^{T}_{i}||q^{T}_{i})],$ (5)
@@
where $KL$ denotes Kullback–Leibler divergence, $q^{I}_{i}=\texttt{softmax}(\{f_{\theta_{v}}(I_{i})\cdot g_{\theta_{l}}(T_{j})\}_{j=1}^{N})$ and $q^{T}_{i}=\texttt{softmax}(\{f_{\theta_{v}}(I_{j})\cdot g_{\theta_{l}}(T_{i})\}_{j=1}^{N})$ are student outputs, and $\tilde{q}^{I}_{j}$ and $\tilde{q}^{T}_{j}$ are teacher outputs which are similarly defined by replacing the student parameter $\theta$ with that of teacher’s $\psi$. Presumably, label smoothing (LS) [55] behaves similarly to what we intended, but we argue that LS would be less effective than EMA SD in terms of mitigating overconfidence issues. See Appendix §B.
@@
We complete the learning objective as a summation of $\mathcal{L}_{\text{MCL-con}}$ and $\mathcal{L}_{\text{SD}}$ with a coefficient $\lambda_{\text{SD}}$, i.e., $\mathcal{L}=\mathcal{L}_{\text{MCL}}+\lambda_{\text{OC}}\mathcal{L}_{\text{OC}}+\lambda_{\text{SD}}\mathcal{L}_{\text{SD}}$. The novel combination of these two components contributes to ensuring a larger smallest singular value of image representation and ID calibration simultaneously, which induces calibrated robust fine-tuning (CaRot) on distribution shifts. Note that this objective function is just one of the possible realizations of our upper bound (Theorem 3.1) on OOD generalization and calibration errors. Further exploration can spawn a more practical algorithm in the future.
@@ 5 Experiments
In § 5.1, we first show empirical evidence of the error bounds that we derived in §3. We then provide experimental setup and main benchmarking results (§5.2) and present further empirical studies (§5.3).
@@ 5.1 Numerical analysis on error bounds
Our theoretical analysis in §3 revealed the possibility of managing OOD classification and calibration errors simultaneously by leveraging a shared quantity over the ID domain (sum of calibration error term and the singular value term). Before conducting real-world evaluations, we verify the theoretical analysis with a toy experiment that simulates distribution shifts. To be specific, we generate a binary classification dataset with 1000-dimensional Gaussian random variables as features where the mean of features are partly shifted across different test environments (ID, OOD). We train a three-layer network with regularization terms: Born-Again-Network (BAN)-style self-distillation [13] for calibration ($\mathcal{L}_{\text{SD}}$) and the orthogonal constraint (shown in eq.(4)) for the singular value ($\mathcal{L}_{\text{OC}}$). Detailed descriptions of the experimental setup are provided in Appendix §A.1.
Figure 3: Analysis of error bounds on synthetic data. Plots on the left side show RHS (x-axis) and LHS (y-axis; MSE for ineq.(12) and ECE for ineq.(11)) of the inequalities in §3. We denote MSE for the mean squared error, $\mathcal{L}_{\text{OC}}$ for the singular value regularization, and $\mathcal{L}_{\text{SD}}$ for the calibration regularization.
@@
Figure 3 visualizes the results of Pearson correlation analysis between the average of $1/\sigma_{min}(\tilde{\Sigma}_{\mathcal{D}_{\text{ID}}})$ and ECE from ID samples and OOD MSE/ECE over 111 trained models. Here, we observe strong correlations between the average of $1/\sigma_{min}(\tilde{\Sigma}_{\mathcal{D}_{\text{ID}}})$ and ID ECE (x-axis), and OOD classification and calibration errors (y-axis). Additional results on the best models per each regularization term are showcased on the Table 5.1, which also indicates that reducing the upper bound results in better OOD generalization and calibration. These analyses demonstrate that Theorem 3.1 empirically holds.
@@ 5.2 Evaluation on distribution shift benchmarks
Benchmark datasets. We consider IN-V2 [49], IN-R [23], IN-A [24], IN-S [60], and ObjectNet [3] as natural shifts of the in-distribution dataset (IN). We refer to the average performance over these five datasets as Avg. Shifts or OOD throughout the following sections unless it is specified as a different dataset, e.g., IN-C [22] which we adopt as a synthetic shift scenario occurred by sensory noises.
Baseline methods. We benchmark CaRot alongside zero-shot inference (ZS) and fine-tuning methods: standard fine-tuning (FT), LP-FT [30], FLYP [17], and Lipsum-FT [42]. Refer §A for further details. In Appendix Table B, C, and D, we compare results with post-hoc robustification method (weight ensemble; WiSE-FT [61]) and post-hoc calibration (temperature scaling; TS [18]) method.
Results on natural shifts. Table 2 and 3 highlight our argument that CaRot significantly enhances both generalization and calibration on OOD data. Under distribution shifts from IN to -V2, -R, -A, -S, and ObjectNet, CaRot favorably compares with the existing best fine-tuning methods by margin of 1.51 and 0.0123 for OOD top-1 accuracy and ECE, respectively, averaged over five shifted datasets. See reliability diagrams in Appendix Figure B for deeper insight on calibration. We further report the results with different backbone models, RN50 and ViT-L/14, in Table 7 (See Table H and I for details). CaRot consistently outperforms the baseline methods for these backbones, too. Furthermore, in Table 6, we provide additional comparisons with CAR-FT [34], Model Stock [26] and ARF [20] [^3] We report the average accuracy over four shifted datasets in Table 6. Results with an asterisk (*) are taken from the original papers. ECE values are not included due to the missing evaluations from the original papers..
Figure 4: IN-C corruption-wise accuracy (top) and ECE (bottom). We evaluate accuracy and ECE over 15 types of image corruption with five corruption severity and report the average performance per corruption. CaRot consistently outperforms baseline methods across diverse corruptions.
Figure 5: Closer look at the effectiveness of CaRot on different corruptions. We provide IN-C accuracy on brightness (left) and elastic transform (right) corruptions. CaRot excels on the coarser corruption such as brightness whereas its effectiveness is weakened on the finer corruption such as elastic transform.
Results on synthetic shifts. In real-world applications, distribution shifts are commonly occurred by sensory noises. To evaluate different fine-tuning methods under such synthetic shifts, we adopt a corrupted version of ImageNet (IN-C) with 15 types of image corruptions over five severities. In Figure 4, we provide corruption-wise accuracy and ECE of each method averaged by five severities. Overall, CaRot consistently outperforms the baseline methods and the actual amount of improvements varying depends on the type of corruptions. Specifically, on the relatively coarser granular corruptions such as Snow, Frost, Fog, Brightness, and Contrast greatly change the semantics of the image (similar to natural shift), CaRot shows remarkably good performance compared to others. Meanwhile, on the finer granular corruptions such as Elastic transform and JPEG compression, the improvements achieved by CaRot become smaller. We present zoom-in results on these two cases in Figure 5.
@@
5.3 Further empirical studies
Ablation study. In Table 4, we provide results of the ablation study to show the impacts of each component of CaRot. In line with our hypothesis, results confirm that all three components boost OOD accuracy and calibration performance. The comparison of adopting and not adopting $\mathcal{L}_{\text{MCL}}$ (we followed the naive fine-tuning approach for the latter) ascertains that employing contrastive loss as a fine-tuning objective is superior to cross-entropy loss for ID/OOD accuracy, consistent with the previous observations [17], and even extends to improvements in calibration as well. The ablations of the orthogonality constraint and adopting self-distillation validate our rationale behind adding the terms to our learning objective, where we expect them to lower the upper bound of OOD classification and calibration errors. Together, constraining the singular values of image representation on MCL and distilling EMA teacher’s predictions show the best results which aligned with results from the demonstration of error bounds in Fig 3. We speculate that learning diverse features by the singular value regularization while being enforced to contribute to reflecting the in-batch similarity structure by EMA SD induces well-restricted solution space [25] otherwise has risk converged to bad solutions (learning diverse features but noise-sensitive). Besides, Table 5 shows the impact of each component by varying the strength coefficients. We observe that the increased intensity of constraint improves ID and OOD performance to some degree, but there is a slight decline when the intensity is too strong. Meanwhile, the strength of self-distillation positively correlated with OOD accuracy and ECE, but there are negative effects on ID accuracy and ECE, which reflects the inevitable trade-off between ID adaptation and OOD generalization (Table G and F in Appendix provide further details).